OpenStudy (anonymous):

f(t)= t^2 +2t, t is greater than or equal to -2, solve f(t)=0

OpenStudy (jhannybean):

\[\large f(t) = t^2 + 2t \ \ , f(t) = 0\] Set the equation equal to 0 \[\large t^2 + 2t =0 \] factor out the like-term t. \[\large t(t+2) = 0\] Solve each one separately \[\large{ t= 0 \\ t+2=0}\] What do you get?